# CAST: Corpus-Aware Self-similarity Enhanced Topic modelling

### v1

@@ 8 Limitations and Future Work
While our proposed model demonstrates superior performance in generating coherent and diverse topics, as well as handling noisy data, there are a few limitations to consider. First, CAST relies on pre-trained embedding models to capture contextual information. These models may not fully represent all possible linguistic nuances and could introduce biases based on their training data. Furthermore, the computational complexity of CAST may be higher compared to traditional topic modelling techniques, particularly for large datasets, due to the additional steps involved in contextualized word embedding generation (see Table 5). Despite these limitations, CAST offers a robust and interpretable approach to topic modelling, leveraging the latest advancements in NLP and neural embeddings.
@@ Appendix B LLM-Based Metrics Design and GPT-4 Validation
As the prompt design is beyond the focus of our paper, we only adapted zero-short prompting. In the future, researchers could adopt more advanced few-short or Chain-of-Thought prompting to achieve a better correlation with human evaluation.
@@ Appendix C Implementation Details
In our experiment, we used UMAP to reduce the dimensionality of the document embeddings and HDBSCAN to cluster the reduced embeddings as suggested in Top2Vec and BERTopic. UMAP has the ability to preserve both local and global structure and be scalable to large datasets (McInnes et al., 2018; Angelov, 2020; Grootendorst, 2022).(McInnes et al., 2018; Angelov, 2020; Grootendorst, 2022). HDBSCAN adopts a minimum spanning tree algorithm based on data point densities to identify clusters and uses a soft-clustering method to identify noise and outlier points.
@@ Appendix D Topic Words Example
Table 8: The top 10 topic words generated by each model on 20NewsGroups with #Topics=10. means the model does not using the self-similarity module to filter words.

### v2

@@ 8 Limitations and Future Work
While our proposed approach demonstrates superior performance in generating coherent and diverse topics, as well as handling noisy data, there are a few limitations to consider. First, CAST relies on pre-trained embedding models to capture contextual information. These models may not fully represent all possible linguistic nuances and could introduce biases based on their training data. Furthermore, the computational complexity of CAST may be higher compared to traditional topic modelling techniques, particularly for large datasets, due to the additional steps involved in contextualized word embedding generation (see Table 5). Despite these limitations, CAST offers a robust and interpretable approach to topic modelling, leveraging the latest advancements in NLP and neural embeddings.
@@ Appendix B LLM-Based Metrics Design and GPT-4 Validation
As the prompt design is beyond the focus of our paper, we only used zero-short prompting. In the future, researchers could use more advanced prompting to achieve a better correlation with human evaluation.
@@ Appendix C Implementation Details
In our experiment, we used UMAP to reduce the dimensionality of the document embeddings and HDBSCAN to cluster the reduced embeddings as suggested in Top2Vec and BERTopic. UMAP has the ability to preserve both local and global structure and be scalable to large datasets (McInnes et al., 2018; Angelov, 2020; Grootendorst, 2022). HDBSCAN adopts a minimum spanning tree algorithm based on data point densities to identify clusters and uses a soft-clustering method to identify noise and outlier points.
@@ Appendix D Topic Words Example
Table 8: The top 10 topic words generated by each model on 20NewsGroups with #Topics=10. means the model does not use the self-similarity module to filter words.